9:00am - Lorraine Walsh 9:45am - China Blue 10:30am - Adrienne Klein 11:15am - Alastair Noble 12:00pm - Doo Eun Choi 12:45pm - Elizabeth Flyntz

1. Lorraine Walsh

• BFA Parsons 84

• MFA Penn 95

• Curation: 4 student shows, 3 shows (2 co-), 2 exhibition design

• Abilities: practicing artist in science/art e.g. “Sound Lines”

• Teaching: Associate Prof (tenured) of New Media at UNC Asheville. “The Night Sky” –Sensing the Radio Sky (student projects, NSF supported)

“nature’s design offers profound potential for reflection on human action. Thus the identification of shared expression in art and science lends itself to cultural reflection, in that these expressions offer exciting vantage points to comment on how societies and individuals interact.” [Problem: will she give up tenure position for 5-year appointment?]

2. China Blue

• BA Cal College of the Arts 92

• MFA Hunter 94

• Curation: 10solo shows + 4 others (3 co-)

• Abilities: grant-writing, marketing, fund-raising, budget Practicing science-artist Extensive community work • Teaching: once as Student Teacher, Hunter (Combined Media); many symposia

and invited addresses. Created “ a new course for high school students on the science of sound in our World.”

3. Adrienne Klein

• BFA Syracuse 73

• MA, MFA SunyAlbany 82,85

• Arts Administration Certificate NYU 99

• Curation: 4 years as college gallery director, plus over 100 events at CUNY grad center

• Abilities: administrative experience in faculty grants NYU

• Teaching: Adjunct Professor in Computer Arts, School of Visual Arts, New York, NY since 2001

“To faithfully characterize the Simons Center and Stony Brook, the gallery should present extraordinary and thought provoking exhibitions. I envision the art program being used to support the dialogues taking place within the Center, and extending the conversation to new audiences.”

4. Alastair Noble

• BA Hull College of Art

• MFA (printmaking) Rutgers

• Curation: 2 shows + 2 at SCGP; 2 more planned

• Abilities: environmental/conceptual artist

• Teaching: 15 years in Art departments; curation?

• Writing: several “Artists Books”

“the gallery would be a laboratory of ideas that demonstrate the synergy between art and science in an open dialogue.”

5. Doo Eun Choi

• BA Hongik (Korea) 99

• MA Hongik 06

• currently doctoral candidate, Plymouth

• Curation: 10 years as museum curator Art Center Nabi, Seoul; many many “curatorial works”

• Presentation “How Science meets Art in 21C” 2006

• Teaching: 3 years teaching Korea, Parsons; digital art, theory, 1 term curating (2005, Dongduck Womens Univ. Seoul)

“As a curator, I treat the space as a laboratory for experimenting with the types of experience and aesthetics that can emerge from combining and recombining scientific knowledge, artistic practices, and historical narratives.”

6. Liz Flyntz

• BA Antioch 02

• MFA Buffalo 12

• Curation: over 12 shows listed

• Writing: several essays, 3 recent

• Abilities: videography, stage manager, marketing

• Teaching: portfolio production, video art analysis 1 course, 1 summer course (?), TA for 1 course.

“My focus is on presenting exhibitions and creative works that bridge science, technology, and contemporary art in order to explore the resonances and dissonances between the fields.” [But not clear from résumé which of the shows accomplish this goal, and how].